Consider the demand for a good. At price Rs 4, the demand for the good is 25 units.

Suppose price of the good increases to Rs 5, and as a result, the demand for the good

falls to 20 units. Calculate the price elasticity?

1
Expert's answer
2021-11-08T17:25:50-0500

Here,P=Rs.4; P"_{1}" ​=Rs.5;


Q=25units ;Q"_{1}" =20 units


Price elasticity of demand ("E_{d}" )=(-)"\frac{P}{Q}\times\frac{\delta Q}{\delta P}"


"\frac{\delta Q}{\delta P}=\frac{20-25}{5-4}=\frac{-5}{1}"


"(E_{d}\u200b )=(-)\\frac{4}{25}\\times\\frac{-5}{1}=0.8"